NAOJ GW Elog Logbook 3.2

Yesterday I uploaded three figures, the first one is the green light power as a function of resistance, since we can only get the resistance from the thermal controller. Today we took smaller step of resistance changing, got better shape of the curve,and finished the convert from the resistance to temperature. The final figure shows in the attachment. We still need to do some modification of the beam to get more symmetric result.

We did the clean-up of the SHG housing today, installed the crystal, mounted the harmonic beam splitter and SHG housing, and finally we got the green light. But during the process of finding the green light, we installed another half-wave plate between the second and third lens of telescope, because of the design of the SHG housing, the polarization is not we thought it should be. In the attachment you can find the situation of the optical bench now. The installation of other two harmonic beam splitter is for the power measurement of the pure green light instead of the mixture of the green and infrared.
We did three groups of measurement:
1.The relationship between the resistance(temperature) and the green light power. Since the thermal controller we are using now cannot show directly the temperature of the SHG, it can only shows the resistance. During this measurement, the diode current of the laser is 1.996A and the temperature of laser crystal is 23.63 degree. From the figure, you can find out that when the resistance is about 3.375, the power reach its peak. According to the manual of the thermal sensor, 3.011 kilo ohm response to 60 degree and 4.147 response to 50 degree. So the phasing matching temperature of the crystal is lower than what we thought it should be at 65 degree.
2.The relationship between the laser crystal temperature and the green light power. Under the circumstance with resistance setting to 3.375 kilo ohm. Since this kind of laser can have two modes in the same time at some temperature which will increase the power of green light to 1.5 times, these temperature points are where we should avoid. The red circle with plus marker in the figure is the temperature we used before. It is in the middle of two unexpected point, so we can still set the temperature there.
3.The last one is to show the linear relationship between the power of infrared and green, since they both linear dependent with diode current of the laser.

Members [Manuel, Matteo L.]
I measured the divergence of the 1310nm laser. Since the beam profiler doesn't work over 1100nm, I used the DET10N-InGaAs Detector to measure the time-dependent intensity while the beam is cut by the chopper. On the oscilloscope, I measure the rising time from 10% to 90% of the gaussian integral and calculate the waist and the divergence. I repeated for several chopper positions along the beam axis.
divergence = 1.0 ± 0.1 mrad.
waist = 420±40 um

The design of the telescope did with some Matlab code, double checked with the JamMT program and it gave out a visually simulation of the telescope, showing in the attachment.During the mounting of the telescope, the maximum power of the Laser has been used. The detail of the practical telescope shows as below:
Focal length
F1--50.2mm
F2--62.9mm
F3--125mm
Distance:(Set the output of EOM as origin)
L1--37.5cm
L2--47.5cm
L3--58.5cm
Final beam waist--82.5cm
Final beam size(radius)--44.0854 micrometer
During the last check, I found out the distance between the Beamsplitter and the first turning mirror is 5cm longer the optical scheme shows, caused by my counting error of the holes.But since behind this part there are some space left, so I thought maybe we can leave it like this.

Today, the machine-shop of the ATC finished the manufacturing of the radiation shield and the window holder which are going to be attached to the inner chamber of our cryostat in the ATC. I took some photographs of the item(s). The next step would be to actually install it in the cryostat but before that I have to order some coated windows to make use of the window holder...

I finished the drawings of the setup of the OpLevs on the optical table for the BS, PR2, and PR3 mirror of KAGRA. They can be found on the JGW document server.

The new laser for the JASMINE scatterometer arrived. However, it seems that we still need some more preparations to use this instrument as no heat sink or ON/OFF switch was delivered together with the laser.

After the BS istallation, I proceed with the implementation of the local controls as it was done for the PR mirror in July.
OPTICAL LEVER
As a first thing, I restored the opltical lever (see first attaced picture). In order to have the PSD voltage sum <15, to avoid saturation, I put a filter in the laser path to reduced the power.
The power reaching the PSD is about 340 mW which corresponds to a voltage sum of 14.8 V.
TRANSFER FUNCTIONS AND CONTROL LOOPS
1) I have measured the mechanical transfer function of the mirror when exciting yaw and pich respectively (plot pag 1 and 2 of the attached document).
2) I have measured the open loop transfer function of the two degree of freedom (plot pag 3).
3) I have closed the two loops and plotted the comparison between the calibrated spectrum when the loops are open and when they are close (plot pag 4 and 5).
CALIBRATION
In order to calibrate the signal I assuemed that the PSD had the same calibration factor measured for the PSD of the same type used for the PR (see entry 276) (maybe this has to be verified..)
I follewed the same procedure explained in entry 276.
n_cal = 0.0071+/- 0.0002 (normalzed calibration)
arm = 0.60 m +/- 0.05 m
V_sum = 14.8 V
Cal_tot = n_cal/(2*arm*V_sum) = (3.04 +/- 0.26 ) e-4 [1/V]

Last friday I observed that the coil driver (#4) used to control the input mirror was not working properly. It also made the crete produce an alarm sound when connected. We have replaced it with another coild driver (#1). In order to do it, we had to restore the connector number 4 that was missing in that coil driver.
We dispose of 6 coil drivers, originally used in TAMA, which are numbered from 1 to 6. (#2 is actually not numbered)
Here a recap on the coil drivers used for filter cavity and input telescoope mirrors. (See also attached pictures)
Coil driver #1: used for the INPUT MIRROR (connector 4 was missing and had to be replaced)
Coil driver #2: used for the BS TELESCOPE (it is actually not numbered)
Coil driver #3: not used (probably is not working properly, to be checked)
Coil driver #4: not used (previously used for the INPUT MIRROR but it deesn't work anymore)
Coil driver #5: used for PR TELESCOPE
Coil driver #6: used for the END MIRROR

Workers: Tatsumi, Takahashi, Yuefan, Eleonora
This morning the BS mirror has been successfully suspended and the suspension has been installed in the BS chamber.
Coils have been connected as shown in the first attached picture.

Today I checked the I/O channals that will be needed for the control of the BS suspension (part of the filter cavity injection systems). Here a brief recap of the CPUs and the I/0 ports used for all the controlled suspensions. See also first picture in the attached file.
INPUT MIRROR (4 input , 4 output)
CPU: SAS_NM1b (133.40.121.74) in central room
INPUT : SC1 MOD 6 /ai 4:7
OUTPUT: DEV1 /ao 0:3
TELESCOPE MIRRORS BS ans PR (4 input, 8 output)
CPU: SAS-NM2b (133.40.121.78) in central room
INPUT: SC1 MOD 3 /ai 0:7 (1, 2 used for PR 5,6 used for BS) NB: #4 doesn't work!
OUTPUT DEV2 /ao 0:7 (1, 2, 3, 4 used for PR 5, 6, 7, 8 used for BS)
END MIRROR (4 input, 4 output)
CPU: SAS-EM2a (133.40.121.75) in end room
INPUT: SC1 MOD3 /ai0:3
OUTPUT: PXI1 Slot3/ao 0:3
NB: Some days ago I had to change the port in the swtch that was connecting the CPUs in the central room to the network. I found out that the port used previusly (see attached picture 2) didn't work anymore and that was the reason why I wasn't able to connect the supervisor PC to the remotes targets.

Workers: Tatsumi, Takahashi, Yuefan, Eleonora
During the installation work, one standoff was dropped off from the mirror.
[Glueing work]
Tatsumi made glueing work for standoff at TAMA.
[Next]
We will make installation work on Thursday afternoon.


[GOOD NEWS] I successfuly hang BS dummy mass with four tungsten wires of 100 micron.
[BAD NEWS] Dumping mass of the intermediate mass is missing.
Together with this mass, strong magnets and two hanging rods are needed.
[21 Nov.] I found these at bottom of rack for absorption measurement system.
Another missing parts are base plates to sit 10 cm diameter masses for hanging JIG.
I recommend to keep whole parts in one container box.

Worker: Yuefan Guo, Daisuke TATSUMI
We failed to set suspension wire of BS mirror.
We try three times and always break the suspesion wire during the pulling up
the aluminum dummy mirror.
I guess that these failure come from systematic reasons.
Therefore, we stop the work today.
-----------------------------------------------------------------------------------------------------------
I know the safy factor of tensile strength is about 2 for 10cm fused silica mirror.
Because BS mirror has different dimensions, I calculate several parameters.
[[ ]]
1) Tensile strength of tungsten wires
2500-3500 N/mm^2
2) Wire dimensions
Diameter 50 micro-meter
Number of wires 4
-->
2500 N/mm^2 * (0.025 * 0.025 * pi) * 4
= 19.6 N
= 2.0 kgf
-------------------------------------------------------------
3) Mass of mirror and dummy
Diameter 1 0cm / Thickness 6 cm Volume = 471 cm^3
Diameter 15 cm / Thickness 4 cm Volume = 707 cm^3
Density of
Fused silica 2.20 g/cm^3
Aluminum alloy 5052 2.68 g/cm^3
Weight of
10cm / Fused silica 1036 g
10cm / Aluminum 1262 g
15cm / Fused silica 1555 g
15cm / Aluminum 1895 g
It means that for 15cm alumimum dummy mirror,
there is almost no safety factor. If tension balance of four wires is not good,
the wires will break easily.
I decide to change the wire diameter to 100 micron.
P.S.
High wire tensions are needed to reduce its mechanical loss.
For stearing mirror of the filter cavity, we do not take car of thermal noises.

Members: Sakai, Manuel
We improved the VI that makes the absorption maps. Now it can also make Circular maps. We added an elapsed/remaining time and progress indicator. We also cleaned up the Front Panel to make it simpler. Next, we will include the filters.

I finished the glueing work for TAMA BS mirror.

I measured the beam height at TAMA center room.
The beam height is about 1200 mm.
This is consistent with Takahashi-san's answer.

I checked some EOM at TAMA.
I found that Yuefan already installed a custom EOM for SHG.
I know the following things.
1) Custom EOM crystal (producted by DELTRONICS Inc.) was assembed into Newport 4003 housing. And then its resonant frequency was tuned to be 15.235 MHz by New Port Inc.
2) To reduce a power density of the passing beam, the crystal has dimensions of 4 x 4 x 40 mm. In addition, holes on the housing were expanded to 4 mm in diameter. Normal New Port EOM 4003 has 2 mm of its aperture.
3) At TAMA site, we have several EOMs as listed bellow.
PRODUCT | Resonant freq. (MHz) | Aperture | |
Custom EOM | 15.235 | 4 mm | already installed for SHG |
NewPort 4003 | 15.24 | 2 mm | |
NewPort 4003 | 76.1 | 2 mm |
